OpenStudy (anonymous):

What is the value of the expression when b = 3?

OpenStudy (anonymous):

@iGreen

OpenStudy (danjs):

4*3^4

OpenStudy (danjs):

4*81

OpenStudy (anonymous):

20,736

OpenStudy (danjs):

324

OpenStudy (anonymous):

4x3 to the fourth

OpenStudy (anonymous):

so 4x3 4 times

OpenStudy (anonymous):

so 12 x 12 x 12 x 12

OpenStudy (anonymous):

no when there is an exponent it means the number times it self that many times

OpenStudy (anonymous):

for example 3 to the 3rd power means 3 x 3 x 3

OpenStudy (anonymous):

so 208

OpenStudy (anonymous):

whats that

OpenStudy (anonymous):

the answer to 12x12x12x12

OpenStudy (anonymous):

no the answer is 20,736 im doing k12 math to dude im 6th grade its easy

OpenStudy (anonymous):

im 8th

OpenStudy (anonymous):

how did you get 208

OpenStudy (danjs):

you must do exponents before multipication, remember your order of operations 4*3^4 = 4*81 = 324

OpenStudy (anonymous):

i added 104 and 104 but i was supposed to multiply

OpenStudy (anonymous):

then how am i doing stuff 2 times as hard as this

OpenStudy (anonymous):

oh crap danjs was right

OpenStudy (anonymous):

because im making my math grade better by going back to redo and do i write the answer like this 20,736

OpenStudy (anonymous):

the answer is 324 give him a medal by clicking best answer

OpenStudy (anonymous):

no its 324

OpenStudy (anonymous):

ok more?

OpenStudy (anonymous):

more questios?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok ask away

OpenStudy (anonymous):

just remember your asking another k12 student thats in 6th grade

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

What is the value of the expression when a = 5 and b = 3?